Suppose P(A)=0.55, P(B)=0.45, P(A∪B')=0.7.

Find P(A∪B). Enter your answer to 3 decimal places.
Answer #1

We know,

P(AU B) = P(A) + P(B) _ P(An B)------------------ 1

but,

P(Acup B{}') = P(A) + P(B{}') - P(Acap B{}')--------------------------2

but

P(Acap B{}') = P(A) - P(Acap B)----------------------3

So therefore put equation 3 in equation 2, We get

P(Acup B{}') = P(A) + P(B{}') - P(A) + P(Acap B)}

P(Acup B{}') = P(B{}') + P(Acap B)

but PlB) = 1-P(B)

We get

P(Acup B{}') = 1 - P(B) + P(Acap B)

So therefore P(Acap B) = P(Acup B{}') - 1 + P(B)

We get P(A n B) = 0.7-1 + 0.45 = 0.15

So therefore Equation 1 becomes

P(AU B) = P(A) + P(B) _ P(An B)

P(A U B) = 0.55 0.45-0.15 = 0.850

P(A U B) = 0.850